OpenStudy (anonymous):

a wire of length 200cm is cut into 2 parts and each part is bent to form a square. if the area of the larger square is 9 times that of the smaller square, find the perimeter of the larger square. (use quadratic formula if possible)

OpenStudy (shinalcantara):

will you please recheck the statement "if the sum of the area..." that line

OpenStudy (zpupster):

area of first square= x^2 area 2nd square = (200-x)^2 first square 9 time bigger so....... 9x^2 = (200-x)^2 9x^2=40000-400x+x^2 add - 9x^2 to both sides 0 = -8x^2 - 400 x + 40000 factor out -8 and divide both sides by -8 then you could use the quadratic formula or just factor

OpenStudy (shinalcantara):